# Security Analysis of ChatGPT: Threats and Privacy Risks

### v1

@@ 1. Introduction
Artificial Intelligence (AI), as a key technology to equip computer systems with learning and reasoning capabilities, has experienced several rounds of development since it was proposed in the mid-20th century. From the early artificial neural networks, the Turing test, and expert systems, to the rise of deep learning and large language models in recent years, AI technology continues to promote the process of social intelligence. In particular, the Transformer model proposed by Google in 2017 laid the foundation for the rapid evolution of Large Language Models (LLMs)(min2023recent, ). Since then, representative models such as BERT, GPT, etc. have demonstrated excellent performance in natural language processing, speech recognition, image generation, and other fields. Currently, major domestic technology enterprises have also launched their large models, such as DeepSeek, Doubao, kimi, etc, to promote the vigorous development of the AI industry.
ChatGPT, as a representative generative AI model launched by OpenAI, has been widely used in many fields such as text writing, code generation, knowledge quizzes, etc., with its powerful language generation capability and good user interaction experience, bringing significant convenience to society. However, at the same time, its risks in terms of cybersecurity, privacy protection, copyright compliance, and educational ethics have gradually emerged (saharia2022photorealistic, ). Studies have shown that ChatGPT may be abused for writing malicious code, generating false information, bypassing security restrictions, etc., while there are privacy issues such as training data leakage, model poisoning, and information reconstruction. Some studies at home and abroad have paid attention to ChatGPT-related attack methods and abusive behaviors, such as prompt injection, phishing email generation, model stealing, etc., and called for strengthening the prudent regulation of its use(takale2024cyber, ).
In this context, this paper centers on the security threats and privacy risks of ChatGPT. Firstly, we sort out the development background of big language models and the technical principles of ChatGPT, and analyze its typical applications in multiple practical scenarios. Subsequently, we summarize and categorize the main security issues and privacy risks faced by ChatGPT, such as malicious content generation, training data leakage, model reverse engineering, etc. Finally, through simulated attack and defense experiments, we validate the feasibility of the relevant attack modes, and explore their potential countermeasures and improvement suggestions(charfeddine2024chatgpt, ).
@@ 2.1. ChatGPT
Since OpenAI first released ChatGPT in 2018, its model architecture and capabilities have gone through several rounds of upgrading and evolution, driving the widespread application of big language models in natural language processing. It was not until the end of 2022 that OpenAI launched the ChatGPT-3.5 version, which truly introduced the dialogue mechanism and was widely promoted in an open form, becoming a key node in the drive to massify generative AI(pan2020privacy, ).
@@
  • ChatGPT-1(2018): ChatGPT-1 is the first version based on the Transformer architecture, adopting a structure containing only a Decoder Block with 12 layers and 117M parameters. The model performs text categorization and generation tasks, and has preliminary Q & A capabilities(ruixue2023insights, ). Through a combination of generative pre-training and discriminative fine-tuning, ChatGPT-1 shows excellent performance in some natural language processing tasks. Nevertheless, due to the lack of conversational capability and limited generalization, it does not yet have the characteristics of a true conversational system.
@@
  • ChatGPT-2 (2019): ChatGPT-2 is significantly extended in architecture with 48 layers and 1542M parameters. The model migrated to a variety of supervised tasks through unsupervised training and achieved leading performance in 7 out of 8 language modeling benchmarks(radford2019language, ). Compared to its predecessor, ChatGPT-2 demonstrates enhanced language understanding and text generation capabilities, enabling simple conversations, story creation, and has even been used to generate content such as phishing emails, showing potential risk of abuse(el2023integration, ).
@@
  • ChatGPT-3 (2020): A leap in scale and capability, with the number of parameters raised to 175B, a 96-layer structure, and a large model based entirely on self-supervised learning. It can accomplish multiple tasks without fine-tuning, which significantly improves the quality of language generation and multi-task adaptation(radford2018improving, ). However, its training process cannot effectively identify harmful information in network data, which may lead to biased or ethical issues in the generated content. In addition, ChatGPT-3 still does not have a true multi-round natural dialog function.
@@
  • ChatGPT-4 (March 2023): ChatGPT-4 is the latest upgraded version at present, and although the parameter scales and architectural details have not been fully disclosed, several evaluations have shown that it significantly outperforms version 3.5 in academic and professional exams(brown2020language, ). Although ChatGPT-4 is mainly offered as a paid service, its improvements in language comprehension, reasoning ability, and response stability make it more suitable for demanding application scenarios.
@@
ChatGPT, as an advanced conversational AI model, has been widely used in customer service, writing, translation, and programming. In customer service scenarios, it can efficiently answer user questions and reduce enterprise labor costs. In the writing and media industry, it can assist in content generation and information refinement, improving creative efficiency. In translation, ChatGPT supports personalized adjustment of tone and style. In programming practice, it can be used for code debugging, annotation, and generation, significantly improving development efficiency(achiam2023gpt, ).
@@ 2.2. Transformer
  • Step 1: As shown in Fig. 2, the model sums the word vectors (word embedding) of each word with its corresponding position vectors (position embedding) in the input stage to obtain the input representation matrix of each word. The word vectors represent the mapping of words into a high-dimensional vector space to capture their semantic features(ashish2017attention, ). In contrast, the position vectors are used to encode information about the order of words in a sentence, compensating for the model’s inability to perceive word order. Where n denotes the number of words in the sentence and d denotes the dimension of the vector. With this embedding approach, the model is able to incorporate positional information while retaining semantic information, thus understanding and processing natural language text more effectively.
@@
  • Step 2: As shown in Fig. 3, the vector matrix of words is input into the Encoder, and after successive layer-by-layer processing of 6 layers of Encoder Blocks, an encoded matrix containing semantic information of all the words in the sentence is finally output.
@@
  • Step 3: As shown in Fig. 4, when the Decoder generates the $(i+1)$ word, it can only attend to the first $i$ previously generated tokens. To enforce this constraint, the Transformer model employs a Masked Multi-Head Attention mechanism, which prevents the model from accessing future positions during decoding(fui2023generative, ). The term ”masked” indicates that the attention vector at each position is computed solely based on the current and previous tokens, thereby avoiding information leakage from subsequent words.
@@ 2.3. Input
Word embeddings are typically pre-trained on large-scale corpora such as Wikipedia, GloVe, Word2Vec, and FastText. These models map each word into a high-dimensional vector space, capturing semantic relationships among words, and providing rich representations for downstream tasks. Unlike RNNs or CNNs, the Transformer model lacks an inherent mechanism for modeling word order(sallam2023chatgpt, ). Therefore, it incorporates Positional Encoding to inject explicit position information into the input representations. Positional encoding vectors have the same dimensionality as word embeddings and vary based on the word’s position in the sentence. In the original Transformer implementation proposed by Vaswani et al., positional encodings are computed using sinusoidal functions as follows in equation 1and equation 2.pospos is the position index of the word in the sentence (starting from 0),$i$ is the dimension index of the positional encoding, and $d$ is the total dimensionality (equal to the word embedding dimension). This method offers two main advantages.After computing the word embedding $x_{i}$ and the positional encoding $p_{i}$, the final input representation for each token is obtained by element-wise addition. The full input matrix for a sentence can be expressed as equation 4. $E$ is the input embedding matrix, $X$ is the word embedding matrix, $P$ is the positional encoding matrix, $n$ is the number of tokens in the sentence, $d$ is the embedding dimension.
@@
    Relative position modeling: Due to the periodic nature of sine and cosine functions, the model can more easily capture relative distances between tokens.
@@ 2.4. Self-Attention Mechanism
As shown in Figure 5, the Transformer model consists of two main components. The Encoder on the left and the Decoder on the right. Within the architecture, there are three Multi-Head Attention (MHA) modules in total, each composed of multiple Self-Attention mechanisms. In the Encoder, there is a single multi-head attention layer. In contrast, the Decoder includes two Multi-Head Attention layers, one of which employs Masked Attention to prevent positions from attending to subsequent positions during training, ensuring the autoregressive property of language generation(huang2020metapoison, ). After each Multi-Head Attention operation, two essential components are applied.
@@ 3.1. Traditional Threats
  • Social Engineering Attacks: Social engineering attacks are a form of attack that exploits human psychological weaknesses to gain access to sensitive information, which is usually accomplished through the stages of information gathering, relationship building, trust manipulation, and malicious payload delivery. Among them, phishing is the most common form, in which attackers lure victims to click on malicious links through forged emails, websites, or SMS to steal identity information or spread malware. Due to the high trust of mobile device users in SMS content, SMS has become one of the high incidence channels for phishing attacks(wang2018stealing, ). With the development of large language models, ChatGPT is used by attackers to generate phishing content with natural and misleading language, which enhances the deception of social engineering attacks. Although OpenAI sets restrictions on certain high-risk keywords, attackers can bypass detection through evasive wording to realize the attack purpose.
@@
  • Malicious Code Generation: Although ChatGPT has built-in content filtering to limit malicious uses, research has shown that attackers can circumvent the restrictions by avoiding prompts or rephrasing requirements to generate malicious code that can be used to steal information. For example, an attacker can use ChatGPT to generate a simple Trojan horse that requires no programming skills. Although the code generated is far less complex and covert than handwritten code by advanced attackers, and often contains logic holes that make it less usable in practice, it significantly lowers the threshold for attack and could pose a potential threat to systems lacking protection mechanisms(krishna2019thieves, ).
@@
  • AI Package Illusion: AI package illusion refers to the fact that when using pre-trained models or open-source AI tools, users have unrealistic expectations about the performance and functionality of the models due to insufficient cognition or blind trust in their capabilities, thus ignoring their scope of application and limitations. This cognitive bias may lead to model performance degradation in inappropriate scenarios, or even output erroneous or misleading results, causing confusion and misjudgment to users. Attackers take advantage of this phenomenon to carry out attacks. For example, when ChatGPT recommends software packages, it may return some non-existent package names due to the hallucination phenomenon. Attackers can disguise malicious code as these ”fictitious packages” and upload them to public code repositories. If other users download and use these malware packages based on ChatGPT’s suggestions, it may lead to system intrusion, data leakage, and other security risks. To minimize such risks, users should rationally evaluate model outputs and avoid relying solely on language model recommendations, especially when it comes to the selection of code, security components, or third-party libraries, which should be manually verified and validated. Figure 12 shows the flow of an attacker propagating a malware package. Figure 13 shows that if a user queries ChatGPT for package suggestions, ChatGPT suggests these malware packages distributed by the attacker, which finally leads to the user using the malware package(carlini2021extracting, ).
@@ 3.2.1. Generation of False Information
Large Language Models often suffer from the phenomenon of ”illusion”, i.e., generating information that seems reasonable but is false, such as misinterpretation of concepts, fabrication of references, etc., which can easily cause misinformation. In 2023, a self-media company used ChatGPT to falsify the ”Gansu Train Collision ”false news to circumvent the platform’s checking and profitability. false news to circumvent platform checking and make a profit, and the related content was disseminated 15,000 times, involving the crime of provoking trouble. This incident reflects the risk of language models being abused by wrongdoers to disseminate false information, which may lead to public opinion misinformation and social harm, and the urgent need to strengthen model governance and content regulation(siau2020artificial, ).
@@ 3.2.2. Data Poisoning
Data poisoning is a typical adversarial method targeting machine learning models, where the core objective is to degrade model performance by manipulating the training data. In a data poisoning attack, the adversary deliberately injects malicious or fabricated samples into the training set, misleading the model to learn incorrect mappings during training, which in turn results in erroneous predictions during inference(lahoti2019ifair, ). Let the model be parameterized by $\theta$ and trained on a dataset $D=D_{\text{clean}}\cup D_{\text{poisoned}}$ with a loss function $\mathcal{L}$. The attacker’s goal is to construct a set of poisoned samples $D_{\text{poisoned}}$ such that the trained model $f_{\theta}$ misclassifies inputs containing specific trigger features—even for previously unseen examples. The objective of such an attack can be formulated as shown in equation 16.
@@
To date, OpenAI has not disclosed the exact sources of its training data. However, it is widely believed that much of the data originates from publicly available internet sources, making the pretraining stage particularly vulnerable to poisoning attacks(courtland2018bias, ). A more critical concern is that current LLMs lack robust and systematic data sanitization mechanisms. Given the high dependency of model performance on data quality, constructing a comprehensive, scalable, and semantically-aware data cleaning system remains an unresolved challenge. Furthermore, adversarial insiders could undermine this process by bypassing verification protocols or manipulating the data curation pipeline, thereby further compromising the robustness of model training.

### v2

@@ 1. Introduction
Artificial Intelligence (AI), as a key technology to equip computer systems with learning and reasoning capabilities, has experienced several rounds of development since it was proposed in the mid-20th century. From the early artificial neural networks, the Turing test, and expert systems, to the rise of deep learning and large language models in recent years, AI technology continues to promote the process of social intelligence(Gallifant et al., 2024). In particular, the Transformer model proposed by Google in 2017 laid the foundation for the rapid evolution of Large Language Models (LLMs)(Min et al., 2023). Since then, representative models such as BERT, GPT, etc. have demonstrated excellent performance in natural language processing, speech recognition, image generation, and other fields. Currently, major domestic technology enterprises have also launched their large models, such as DeepSeek, Doubao, kimi, etc, to promote the vigorous development of the AI industry.
ChatGPT, as a representative generative AI model launched by OpenAI, has been widely used in many fields such as text writing, code generation, knowledge quizzes, etc., with its powerful language generation capability and good user interaction experience, bringing significant convenience to society. However, at the same time, its risks in terms of cybersecurity, privacy protection, copyright compliance, and educational ethics have gradually emerged (Saharia et al., 2022). Studies have shown that ChatGPT may be abused for writing malicious code, generating false information, bypassing security restrictions, etc., while there are privacy issues such as training data leakage, model poisoning, and information reconstruction. Some studies at home and abroad have paid attention to ChatGPT-related attack methods and abusive behaviors, such as prompt injection, phishing email generation, model stealing, etc., and called for strengthening the prudent regulation of its use(Takale et al., 2024).
In this context, this paper centers on the security threats and privacy risks of ChatGPT. Firstly, we sort out the development background of big language models and the technical principles of ChatGPT, and analyze its typical applications in multiple practical scenarios. Subsequently, we summarize and categorize the main security issues and privacy risks faced by ChatGPT, such as malicious content generation, training data leakage, model reverse engineering, etc. Finally, through simulated attack and defense experiments, we validate the feasibility of the relevant attack modes, and explore their potential countermeasures and improvement suggestions(Charfeddine et al., 2024).
@@ 2.1. ChatGPT
Since OpenAI first released ChatGPT in 2018, its model architecture and capabilities have gone through several rounds of upgrading and evolution, driving the widespread application of big language models in natural language processing(Lucchi, 2024). It was not until the end of 2022 that OpenAI launched the ChatGPT-3.5 version, which truly introduced the dialogue mechanism and was widely promoted in an open form, becoming a key node in the drive to massify generative AI(Pan et al., 2020).
@@
  • ChatGPT-1(2018): ChatGPT-1 is the first version based on the Transformer architecture, adopting a structure containing only a Decoder Block with 12 layers and 117M parameters. The model performs text categorization and generation tasks, and has preliminary Q & A capabilities(Ruixue et al., 2023). Through a combination of generative pre-training and discriminative fine-tuning, ChatGPT-1 shows excellent performance in some natural language processing tasks. Nevertheless, due to the lack of conversational capability and limited generalization, it does not yet have the characteristics of a true conversational system.
@@
  • ChatGPT-2 (2019): ChatGPT-2 is significantly extended in architecture with 48 layers and 1542M parameters. The model migrated to a variety of supervised tasks through unsupervised training and achieved leading performance in 7 out of 8 language modeling benchmarks(Radford et al., 2019). Compared to its predecessor, ChatGPT-2 demonstrates enhanced language understanding and text generation capabilities, enabling simple conversations, story creation, and has even been used to generate content such as phishing emails, showing potential risk of abuse(El Saddik and Ghaboura, 2023).
@@
  • ChatGPT-3 (2020): A leap in scale and capability, with the number of parameters raised to 175B, a 96-layer structure, and a large model based entirely on self-supervised learning. It can accomplish multiple tasks without fine-tuning, which significantly improves the quality of language generation and multi-task adaptation(Radford et al., 2018). However, its training process cannot effectively identify harmful information in network data, which may lead to biased or ethical issues in the generated content. In addition, ChatGPT-3 still does not have a true multi-round natural dialog function(Adhikari et al., 2024).
@@
  • ChatGPT-4 (March 2023): ChatGPT-4 is the latest upgraded version at present, and although the parameter scales and architectural details have not been fully disclosed, several evaluations have shown that it significantly outperforms version 3.5 in academic and professional exams(Brown et al., 2020). Although ChatGPT-4 is mainly offered as a paid service, its improvements in language comprehension, reasoning ability, and response stability make it more suitable for demanding application scenarios(Abdaljaleel et al., 2024).
@@
ChatGPT, as an advanced conversational AI model, has been widely used in customer service, writing, translation, and programming. In customer service scenarios, it can efficiently answer user questions and reduce enterprise labor costs(Briganti, 2024). In the writing and media industry, it can assist in content generation and information refinement, improving creative efficiency. In translation, ChatGPT supports personalized adjustment of tone and style. In programming practice, it can be used for code debugging, annotation, and generation, significantly improving development efficiency(Achiam et al., 2023).
@@ 2.2. Transformer
  • Step 1: As shown in Fig. 2, the model sums the word vectors (word embedding) of each word with its corresponding position vectors (position embedding) in the input stage to obtain the input representation matrix of each word. The word vectors represent the mapping of words into a high-dimensional vector space to capture their semantic features(Ashish, 2017). In contrast, the position vectors are used to encode information about the order of words in a sentence, compensating for the model’s inability to perceive word order. Where n denotes the number of words in the sentence and d denotes the dimension of the vector. With this embedding approach, the model is able to incorporate positional information while retaining semantic information, thus understanding and processing natural language text more effectively(Jiang et al., 2024).
@@
  • Step 2: As shown in Fig. 3, the vector matrix of words is input into the Encoder, and after successive layer-by-layer processing of 6 layers of Encoder Blocks, an encoded matrix containing semantic information of all the words in the sentence is finally output(Zhao et al., 2024).
@@
  • Step 3: As shown in Fig. 4, when the Decoder generates the $(i+1)$ word, it can only attend to the first $i$ previously generated tokens. To enforce this constraint, the Transformer model employs a Masked Multi-Head Attention mechanism, which prevents the model from accessing future positions during decoding(Fui-Hoon Nah et al., 2023). The term ”masked” indicates that the attention vector at each position is computed solely based on the current and previous tokens, thereby avoiding information leakage from subsequent words(Tang et al., 2024).
@@ 2.3. Input
Word embeddings are typically pre-trained on large-scale corpora such as Wikipedia, GloVe, Word2Vec, and FastText. These models map each word into a high-dimensional vector space, capturing semantic relationships among words, and providing rich representations for downstream tasks. Unlike RNNs or CNNs, the Transformer model lacks an inherent mechanism for modeling word order(Sallam, 2023). Therefore, it incorporates Positional Encoding to inject explicit position information into the input representations. Positional encoding vectors have the same dimensionality as word embeddings and vary based on the word’s position in the sentence(Li et al., 2024a). In the original Transformer implementation proposed by Vaswani et al., positional encodings are computed using sinusoidal functions as follows in equation 1and equation 2.pospos is the position index of the word in the sentence (starting from 0),$i$ is the dimension index of the positional encoding, and $d$ is the total dimensionality (equal to the word embedding dimension). This method offers two main advantages.After computing the word embedding $x_{i}$ and the positional encoding $p_{i}$, the final input representation for each token is obtained by element-wise addition. The full input matrix for a sentence can be expressed as equation 4. $E$ is the input embedding matrix, $X$ is the word embedding matrix, $P$ is the positional encoding matrix, $n$ is the number of tokens in the sentence, $d$ is the embedding dimension(Hou et al., 2024).
@@
    Relative position modeling: Due to the periodic nature of sine and cosine functions, the model can more easily capture relative distances between tokens(Hatamizadeh and Kautz, 2025).
@@ 2.4. Self-Attention Mechanism
As shown in Figure 5, the Transformer model consists of two main components. The Encoder on the left and the Decoder on the right. Within the architecture, there are three Multi-Head Attention (MHA) modules in total, each composed of multiple Self-Attention mechanisms(Zheng et al., 2024). In the Encoder, there is a single multi-head attention layer. In contrast, the Decoder includes two Multi-Head Attention layers, one of which employs Masked Attention to prevent positions from attending to subsequent positions during training, ensuring the autoregressive property of language generation(Huang et al., 2020). After each Multi-Head Attention operation, two essential components are applied.
@@ 3.1. Traditional Threats
  • Social Engineering Attacks: Social engineering attacks are a form of attack that exploits human psychological weaknesses to gain access to sensitive information, which is usually accomplished through the stages of information gathering, relationship building, trust manipulation, and malicious payload delivery. Among them, phishing is the most common form, in which attackers lure victims to click on malicious links through forged emails, websites, or SMS to steal identity information or spread malware. Due to the high trust of mobile device users in SMS content, SMS has become one of the high incidence channels for phishing attacks(Wang and Gong, 2018). With the development of large language models, ChatGPT is used by attackers to generate phishing content with natural and misleading language, which enhances the deception of social engineering attacks. Although OpenAI sets restrictions on certain high-risk keywords, attackers can bypass detection through evasive wording to realize the attack purpose(Birthriya et al., 2025).
@@
  • Malicious Code Generation: Although ChatGPT has built-in content filtering to limit malicious uses, research has shown that attackers can circumvent the restrictions by avoiding prompts or rephrasing requirements to generate malicious code that can be used to steal information(Zhang et al., 2024). For example, an attacker can use ChatGPT to generate a simple Trojan horse that requires no programming skills. Although the code generated is far less complex and covert than handwritten code by advanced attackers, and often contains logic holes that make it less usable in practice, it significantly lowers the threshold for attack and could pose a potential threat to systems lacking protection mechanisms(Krishna et al., 2019).
@@
  • AI Package Illusion: AI package illusion refers to the fact that when using pre-trained models or open-source AI tools, users have unrealistic expectations about the performance and functionality of the models due to insufficient cognition or blind trust in their capabilities, thus ignoring their scope of application and limitations. This cognitive bias may lead to model performance degradation in inappropriate scenarios, or even output erroneous or misleading results, causing confusion and misjudgment to users. Attackers take advantage of this phenomenon to carry out attacks(Leong, 2024). For example, when ChatGPT recommends software packages, it may return some non-existent package names due to the hallucination phenomenon. Attackers can disguise malicious code as these ”fictitious packages” and upload them to public code repositories. If other users download and use these malware packages based on ChatGPT’s suggestions, it may lead to system intrusion, data leakage, and other security risks. To minimize such risks, users should rationally evaluate model outputs and avoid relying solely on language model recommendations, especially when it comes to the selection of code, security components, or third-party libraries, which should be manually verified and validated. Figure 12 shows the flow of an attacker propagating a malware package. Figure 13 shows that if a user queries ChatGPT for package suggestions, ChatGPT suggests these malware packages distributed by the attacker, which finally leads to the user using the malware package(Carlini et al., 2021).
@@ 3.2.1. Generation of False Information
Large Language Models often suffer from the phenomenon of ”illusion”, i.e., generating information that seems reasonable but is false, such as misinterpretation of concepts, fabrication of references, etc., which can easily cause misinformation. In 2023, a self-media company used ChatGPT to falsify the ”Gansu Train Collision ”false news to circumvent the platform’s checking and profitability. false news to circumvent platform checking and make a profit, and the related content was disseminated 15,000 times, involving the crime of provoking trouble. This incident reflects the risk of language models being abused by wrongdoers to disseminate false information, which may lead to public opinion misinformation and social harm, and the urgent need to strengthen model governance and content regulation(Siau and Wang, 2020).
@@ 3.2.2. Data Poisoning
Data poisoning is a typical adversarial method targeting machine learning models, where the core objective is to degrade model performance by manipulating the training data. In a data poisoning attack, the adversary deliberately injects malicious or fabricated samples into the training set, misleading the model to learn incorrect mappings during training, which in turn results in erroneous predictions during inference(Lahoti et al., 2019). Let the model be parameterized by $\theta$ and trained on a dataset $D=D_{\text{clean}}\cup D_{\text{poisoned}}$ with a loss function $\mathcal{L}$. The attacker’s goal is to construct a set of poisoned samples $D_{\text{poisoned}}$ such that the trained model $f_{\theta}$ misclassifies inputs containing specific trigger features—even for previously unseen examples. The objective of such an attack can be formulated as shown in equation 16.
@@
To date, OpenAI has not disclosed the exact sources of its training data. However, it is widely believed that much of the data originates from publicly available internet sources, making the pretraining stage particularly vulnerable to poisoning attacks(Courtland, 2018). A more critical concern is that current LLMs lack robust and systematic data sanitization mechanisms. Given the high dependency of model performance on data quality, constructing a comprehensive, scalable, and semantically-aware data cleaning system remains an unresolved challenge. Furthermore, adversarial insiders could undermine this process by bypassing verification protocols or manipulating the data curation pipeline, thereby further compromising the robustness of model training(He et al., 2025).